# Preconditioning without a preconditioner: faster ridge-regression and Gaussian sampling with randomized block Krylov subspace methods

### v1

@@ Abstract
We describe a randomized variant of the block conjugate gradient method for solving a single positive-definite linear system of equations. Our method provably outperforms preconditioned conjugate gradient with a broad-class of Nyström-based preconditioners, without ever explicitly constructing a preconditioner. In analyzing our algorithm, we derive theoretical guarantees for new variants of Nyström preconditioned conjugate gradient which may be of separate interest. We also describe how our approach yields state-of-the-art algorithms for key data-science tasks such as computing the entire ridge regression regularization path and generating multiple independent samples from a high-dimensional Gaussian distribution.
Machine Learning, ICML
@@ 1 Introduction
$\mathbf{A}_{\mu}\mathbf{x}=\mathbf{b},\quad\mathbf{A}_{\mu}:=\mathbf{A}+\mu\mathbf{I},$ (1)
where $\mathbf{A}\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\mu\geq 0$ is a critical task across the computational sciences. Systems of the form Equation 1 arise in a variety of settings, but we are particularly motivated by two tasks in machine learning and data science: solving ridge-regression problems and sampling Gaussian vectors.
Ridge regression. Given a data matrix $\mathbf{Z}\in\mathbb{R}^{n\times d}$, a source term $\mathbf{f}\in\mathbb{R}^{n}$, and a regularization parameter $\mu\geq 0$, the ridge regression problem [^1] When $\mu=0$, Equation 2 is just standard linear least squares. is
$\min_{\mathbf{x}\in\mathbb{R}^{d}}\|\mathbf{Z}\mathbf{x}-\mathbf{f}\|^{2}+\mu\|\mathbf{x}\|^{2}.$ (2)
A direct computation shows that the solution to Equation 2 is also the solution to
$(\mathbf{Z}^{\mathsf{T}}\mathbf{Z}+\mu\mathbf{I})\mathbf{x}=\mathbf{Z}^{\mathsf{T}}\mathbf{f};$ (3)
i.e. a system of the form Equation 1 with $\mathbf{A}=\mathbf{Z}^{\mathsf{T}}\mathbf{Z}$ and $\mathbf{b}=\mathbf{Z}^{\mathsf{T}}\mathbf{f}$. In a number of applications we are interested in the whole regularization path; i.e. $\mathbf{x}=\mathbf{x}_{\mu}$ for all values $\mu\geq 0$; e.g. in order to perform cross-validation to select a value of $\mu$ to use for future predictions.
@@
$\mathbf{A}^{1/2}\mathbf{b}=\frac{2}{\pi}\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}\,\mathrm{d}z.$ (4)
The term $(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}$ in the integrand of Equation 4 is of the form Equation 1 with $\mu=z^{2}$. Often we are interested in sampling several Gaussian vectors with the same covariance matrix.
Figure 1: Comparison of the convergence (in terms of matrix-loads) of our proposed randomized variant of the block conjugate gradient method (Refer to caption) against standard conjugate gradient (Refer to caption) and the state of the art Nyström preconditioned conjugate gradient (Frangella et al., 2023) with various choices of hyperparameters (Refer to caption). Our method outperforms these existing methods without the need for selecting hyperparameters, which may be difficult to do effectively in practice (see Theorem 3.2).
@@ 1.1 Computational assumptions
Throughout this paper, we assume $\mathbf{A}$ is accessed only through matrix-vector products $\mathbf{x}\mapsto\mathbf{A}\mathbf{x}$ or block matrix-vector products $[\mathbf{x}_{1}~{}\cdots~{}\mathbf{x}_{m}]\mapsto[\mathbf{A}\mathbf{x}_{1}~{}\cdots~{}\mathbf{A}\mathbf{x}_{m}]$. While one can simulate a block matrix-vector product using $m$ matrix-vector products, the parallelism inherent in modern computing architectures means the cost of block products is often nearly independent of the block size $m$ (so long as $m$ is not too large). Allowing for more efficient data-access is widely understood as one of the major benefits of many randomized linear algebra algorithms (Halko et al., 2011; Li et al., 2017; Tropp & Webber, 2023).
The algorithms we propose in this paper are designed to economize the number of times $\mathbf{A}$ is loaded into memory (henceforth referred to as matrix-loads), and not other costs such as matrix-vector products or storage. In settings where matrix-loads are not the dominant cost, the proposed algorithms may not yield significant advantages. The costs of the algorithm are discussed more in Section 3.2, and we explore them in the numerical experiments in Section 5.
@@ 1.2 Notation
We denote the eigenvalues of $\mathbf{A}$ by $\lambda_{1}\geq\cdots\geq\lambda_{d}$ and of those of an arbitrary positive definite matrix $\mathbf{M}$ by $\lambda_{1}(\mathbf{M})\geq\cdots\geq\lambda_{d}(\mathbf{M})$. The condition number of $\mathbf{M}$ is $\kappa(\mathbf{M}):=\lambda_{1}(\mathbf{M})/\lambda_{d}(\mathbf{M})$ and the spectral norm is $\|\mathbf{M}\|:=\lambda_{1}(\mathbf{B})$. The $\mathbf{M}$-norm of a vector $\mathbf{x}$ is $\|\mathbf{x}\|_{\mathbf{M}}:=\|\mathbf{M}^{1/2}\mathbf{x}\|=\sqrt{\mathbf{x}^{\mathsf{T}}\mathbf{M}\mathbf{x}}$.
@@ 1.3 Organization
In Section 2 we describe key background on Krylov subspace methods and preconditioning. Our algorithm and main theoretical results are described in Section 3 and proved in Section 4. Numerical experiments are provided in Section 5. Finally show how the theory we develop can be used to sample Gaussian vectors in Section 6.
@@ 2 Background
Krylov Subspace Methods (KSM) make use of the Krylov subspace [^2] Note that $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b})=\mathcal{K}_{t}(\mathbf{A},\mathbf{b})$, for any scalar $\mu$.
$\mathcal{K}_{t}(\mathbf{A},\mathbf{b}):=\operatorname{span}\{\mathbf{b},\mathbf{A}\mathbf{b},\ldots,\mathbf{A}^{t-1}\mathbf{b}\},$ (5)
and are perhaps the most widely used algorithms for solving systems of the form Equation 1. The methods we discuss in this section are standard, and for mathematical simplicity, we define them by their optimality conditions rather than algorithmically. We discuss implementation details in Appendix A; see also (Greenbaum, 1997; Saad, 2003; Meurant, 2006) for comprehensive treatments of these methods.
@@ 2.1 Preconditioned Conjugate Gradient
The preconditioned conjugate gradient algorithm (PCG) is perhaps the most powerful KSM for positive edfinite linear systems. PCG makes use of a positive definite preconditioner $\mathbf{P}_{\mu}$ that transforms the system $\mathbf{A}_{\mu}\mathbf{x}=\mathbf{b}$ into the system
$(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\mathbf{y}=\mathbf{P}_{\mu}^{-1/2}\mathbf{b},\quad\mathbf{P}_{\mu}^{1/2}\mathbf{x}=\mathbf{y},$ (6)
@@
$\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b};\mathbf{P}_{\mu}):=\mathcal{K}_{t}(\mathbf{P}_{\mu}^{-1}\mathbf{A},\mathbf{P}_{\mu}^{-1}\mathbf{b}).$ (7)
@@
The iterate $\mathsf{pcg}_{t}(\mu)$ can be computed using $t-1$ matrix-vector products with $\mathbf{A}$ (and $t$ products with $\mathbf{P}_{\mu}^{-1}$). We discuss implementation details in Appendix A.
The simplest choice of preconditioner is $\mathbf{P}_{\mu}=\mathbf{I}$, which yields the standard conjugate gradient algorithm (CG) (Hestenes & Stiefel, 1952).
@@ 2.1.1 Convergence
PCG (and hence CG) satisfy a well-known convergence guarantee in terms of the condition number of the preconditioned system $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$; see e.g. (Greenbaum, 1997).
@@
Let $\mathbf{P}_{\mu}$ be any preconditioner and let
$\varepsilon_{t}(\mu):=2\exp\left(-\frac{2t}{\sqrt{\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})}}\right).$
Then the $t$-th preconditioned-CG iterate corresponding to the preconditioner $\mathbf{P}_{\mu}$ satisfies
$\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{pcg}_{t}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq\varepsilon_{t}(\mu).$
Corollary 2.3 implies that if we can find $\mathbf{P}_{\mu}$ so that $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$ is small, then PCG converges rapidly. The choice of $\mathbf{P}_{\mu}$ minimizing this condition number is of course $\mathbf{P}_{\mu}=\mathbf{A}_{\mu}^{-1}$, but this is not a practical preconditioner; if we knew $\mathbf{A}_{\mu}^{-1}$ we could easily compute the solution $\mathbf{A}_{\mu}^{-1}\mathbf{b}$. Thus, finding a suitable choice of $\mathbf{P}_{\mu}$, which balances improvements to the convergence of PCG with the cost of building/applying $\mathbf{P}_{\mu}$ is critical.
@@ 2.2 Deflation preconditioner
If $\mathbf{A}$ is poorly conditioned due to the presence of $r$ eigenvalues much larger than the remaining $n-r$ eigenvalues, then we might hope to learn a good approximation of the top $r$ eigenvalues and “correct” this ill-conditioning. Towards this end, suppose we have a good rank-$r$ approximation $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ to $\mathbf{A}$, where $\mathbf{U}$ has orthonormal columns and $\mathbf{D}$ is diagonal. Intuitively, this low-rank approximation contains the information needed to remove the top eigenvalues of $\mathbf{A}$, thereby reducing the condition number of the preconditioned system. In particular, one can form the preconditioner
$\mathbf{P}_{\mu}:=\frac{1}{\theta+\mu}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$ (8)
and it is not hard to verify that
$\mathbf{P}_{\mu}^{-1}=(\theta+\mu)\mathbf{U}(\mathbf{D}+\mu\mathbf{I})^{-1}\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}).$ (9)
Here $\theta>0$ is a parameter that must be chosen along with the factorization $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$.
Applying $\mathbf{P}_{\mu}^{-1}$ to a vector requires $O(dr)$ arithmetic operations, which is relatively cheap if $r\ll d$. Hence, so long as a reasonable factorization $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ can be obtained, Equation 8 can be used as a preconditioner; see (Frank & Vuik, 2001; Gutknecht, 2012) and the references within.

### v2

# Preconditioning without a preconditioner using randomized block Krylov subspace methods

@@ Abstract
We describe a randomized variant of the block conjugate gradient method for solving a single positive-definite linear system of equations. This method provably outperforms preconditioned conjugate gradient with a broad-class of Nyström-based preconditioners, without ever explicitly constructing a preconditioner. In analyzing our algorithm, we derive theoretical guarantees for new variants of Nyström preconditioned conjugate gradient which may be of separate interest. We also describe how our approach yields fast algorithms for key data-science tasks such as computing the entire ridge regression regularization path and generating multiple independent samples from a high-dimensional Gaussian distribution.
keywords:
preconditioning, conjugate gradient, block Krylov subspace methods
{MSCcodes}
65F08, 65F60, 65F50, 68Q25
@@ 1 Introduction
(1) $\mathbf{A}_{\mu}\mathbf{x}=\mathbf{b},\quad\mathbf{A}_{\mu}:=\mathbf{A}+\mu\mathbf{I},$
where $\mathbf{A}\in\mathbb{R}^{d\times d}$ is symmetric positive definite and $\mu\geq 0$ is a critical task across the computational sciences. Systems of the form Eq. 1 arise in a variety of settings, including the following:
Positive definite linear systems. When $\mu=0$, Eq. 1 is simply the task of solving a positive definite linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$, one of the most common problems in numerical linear algebra.
Ridge regression. Given a data matrix $\mathbf{Z}\in\mathbb{R}^{n\times d}$, a source term $\mathbf{f}\in\mathbb{R}^{n}$, and a regularization parameter $\mu\geq 0$, the ridge regression problem is
(2) $\min_{\mathbf{x}\in\mathbb{R}^{d}}\|\mathbf{Z}\mathbf{x}-\mathbf{f}\|^{2}+\mu\|\mathbf{x}\|^{2}.$
A direct computation shows that the solution to Eq. 2 is also the solution to
(3) $(\mathbf{Z}^{\mathsf{T}}\mathbf{Z}+\mu\mathbf{I})\mathbf{x}=\mathbf{Z}^{\mathsf{T}}\mathbf{f};$
i.e. a system of the form Eq. 1 with $\mathbf{A}=\mathbf{Z}^{\mathsf{T}}\mathbf{Z}$ and $\mathbf{b}=\mathbf{Z}^{\mathsf{T}}\mathbf{f}$. In a number of applications we are interested in the whole regularization path; i.e. $\mathbf{x}=\mathbf{x}_{\mu}$ for all values $\mu\geq 0$; e.g. in order to perform cross-validation to select a value of $\mu$ to use for future predictions.
@@
(4) $\mathbf{A}^{1/2}\mathbf{b}=\frac{2}{\pi}\int_{0}^{\infty}\mathbf{A}(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}\,\mathrm{d}z.$
The term $(\mathbf{A}+z^{2}\mathbf{I})^{-1}\mathbf{b}$ in the integrand of Eq. 4 is of the form Eq. 1 with $\mu=z^{2}$. Often we are interested in sampling several Gaussian vectors with the same covariance matrix.
Figure 1: Relative error $\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{alg}\|_{\mathbf{A}}/\|\mathbf{A}^{-1}\mathbf{b}\|_{\mathbf{A}}$ in terms of matrix-loads (left) and wall-clock time (right) for our proposed randomized variant of the block conjugate gradient method (Refer to caption), standard conjugate gradient (Refer to caption), Nyström preconditioned conjugate gradient from [23] (Refer to caption), and generalizations Nyström preconditioned conjugate gradient using higher-depth Nyström approximations (Refer to caption). Our method outperforms all these methods without the need for selecting hyperparameters (see Theorem 3.4), which may be difficult to do effectively in practice. In this example, the runtime is dominated by the cost of loading chunks of the matrix into memory, so the wall-clock-time is nearly proportional to matrix-loads. Full experiment description in Section 7.1.
1.1 (Block) Krylov subspace methods
Central to this paper are a class of algorithms for solving systems of the form Eq. 1 called Krylov Subspace Methods (KSMs). KSMs make use of the so-called Krylov subspace
(5) $\mathcal{K}_{t}(\mathbf{A},\mathbf{b}):=\operatorname{span}\{\mathbf{b},\mathbf{A}\mathbf{b},\ldots,\mathbf{A}^{t-1}\mathbf{b}\},$
which can be efficiently computed using matrix-vector products with $\mathbf{A}$. Since it is often of interest to solve Eq. 1 for multiple values of $\mu$, we repeatedly make use of the fact that the Krylov subspace is shift-invariant; i.e. that for any scalar $\mu$, $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b})=\mathcal{K}_{t}(\mathbf{A},\mathbf{b})$.
In this paper, we advocate for the use of generalization of KSMs called block-KSMs. Given a matrix $\mathbf{B}\in\mathbb{R}^{d\times m}$ (typically $m\ll d$) with columns $\mathbf{b}^{(1)},\ldots,\mathbf{b}^{(m)}$, the block Krylov subspace is defined as
(6) $\mathcal{K}_{t}(\mathbf{A},\mathbf{B}):=\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(1)})+\cdots+\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(m)}).$
That is, $\mathcal{K}_{t}(\mathbf{A},\mathbf{B})$ is the space consisting of all linear combinations of vectors in the Krylov subspaces $\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(1)}),\ldots,\mathcal{K}_{t}(\mathbf{A},\mathbf{b}^{(m)})$.
As we discuss in Section 1.2, compared to their single-vector counterparts, block-KSMs have a number of computational benefits on modern computational architectures. As such, block-KSMs are widely used for tasks in numerical linear algebra such as low-rank approximation and eigenvalue approximation, where they are known to satisfy strong theoretical guarantees [27, 55, 34, 45, 38, 40, 61].
Interestingly, the use of block-KSMs for solving problems like Eq. 1 is more limited (see Section 5 for a discussion), and single-vector methods such as the conjugate gradient algorithm are more popular. A high-level aim of this paper is to encourage further study on the use of block-KSMs for tasks such as Eq. 1. In particular, we believe that techniques from randomized numerical linear algebra have the potential to provide new insights and theoretical guarantees for block-KSMs used to solve Eq. 1.
@@ 1.2 Computational assumptions
Throughout this paper, we assume $\mathbf{A}$ is accessed only through matrix-vector products $\mathbf{x}\mapsto\mathbf{A}\mathbf{x}$ or block matrix-vector products $\mathbf{x}_{1}~\cdots~\mathbf{x}_{m}]\mapsto[\mathbf{A}\mathbf{x}_{1}~\cdots~\mathbf{A}\mathbf{x}_{m}]$. While one can simulate a block matrix-vector product using $m$ matrix-vector products, in many settings (see e.g. Fig. 1) the cost of block products is often nearly independent of the block size $m$ (so long as $m$ is not too large). Allowing for more efficient data-access is widely understood as one of the major benefits of many randomized linear algebra algorithms [40, 61].
The algorithms we propose in this paper are designed to economize the number of times $\mathbf{A}$ is loaded into memory (henceforth referred to as matrix-loads), and not other costs such as matrix-vector products, floating point operations, or storage. In settings where matrix-loads are not the dominant cost, the proposed algorithms may not yield significant advantages. The costs of the algorithm from this paper are discussed more in Section 3.2, and we explore them in the numerical experiments in Section 7. A detailed understanding of tradeoffs in various costs is somewhat beyond the scope of this paper, but is of practical importance.
@@ 1.3 Notation
We denote the eigenvalues of $\mathbf{A}$ by $\lambda_{1}\geq\cdots\geq\lambda_{d}$. We write $\kappa=\lambda_{1}/\lambda_{d}$ and $\kappa_{r+1}(\mu):=(\lambda_{r+1}+\mu)/(\lambda_{d}+\mu)$, as the condition number of $\mathbf{A}_{\mu}$ with the top-$r$ eigenvalues removed. The eigenvalues of an arbitrary positive definite matrix $\mathbf{M}$ are $\lambda_{1}(\mathbf{M})\geq\cdots\geq\lambda_{d}(\mathbf{M})$. The condition number of $\mathbf{M}$ is $\kappa(\mathbf{M}):=\lambda_{1}(\mathbf{M})/\lambda_{d}(\mathbf{M})$ and the spectral norm is $\|\mathbf{M}\|:=\lambda_{1}(\mathbf{B})$. The $\mathbf{M}$-norm of a vector $\mathbf{x}$ is $\|\mathbf{x}\|_{\mathbf{M}}:=\|\mathbf{M}^{1/2}\mathbf{x}\|=\sqrt{\mathbf{x}^{\mathsf{T}}\mathbf{M}\mathbf{x}}$.
@@ 1.4 Organization
In Section 2 we describe key background on Krylov subspace methods and preconditioning. Our algorithm and main conceptual results are described in Section 3. Explicit and simple to use probabilistic results are proved in Section 4. We discuss our bounds in the context of past work in Section 5. Numerical experiments are provided in Section 7. Finally, we show how the theory we develop can be used to sample Gaussian vectors in Section 6.
@@ 2 Background
In this section we provide an overview of a number of relevant algorithms. The methods we discuss in this section are standard, and for mathematical simplicity, we define them by their optimality conditions rather than algorithmically (all of our bounds assume exact arithmetic); eee [29, 56, 42, etc.] for comprehensive treatments of these methods.
@@ 2.1 Preconditioned Conjugate Gradient
The preconditioned conjugate gradient algorithm (PCG) is a powerful KSM for solving positive definite linear systems. PCG makes use of a positive definite preconditioner $\mathbf{P}_{\mu}$ that transforms the system $\mathbf{A}_{\mu}\mathbf{x}=\mathbf{b}$ into the system
(7) $(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})\mathbf{y}=\mathbf{P}_{\mu}^{-1/2}\mathbf{b},\quad\mathbf{P}_{\mu}^{1/2}\mathbf{x}=\mathbf{y},$
@@
(8) $\mathcal{K}_{t}(\mathbf{A}_{\mu},\mathbf{b};\mathbf{P}_{\mu}):=\mathcal{K}_{t}(\mathbf{P}_{\mu}^{-1}\mathbf{A},\mathbf{P}_{\mu}^{-1}\mathbf{b}).$
@@
The iterate $\mathsf{pcg}_{t}(\mu)$ can be computed using t1t-1 matrix-vector products with $\mathbf{A}$ (and $t$ products with $\mathbf{P}_{\mu}^{-1}$).
The simplest choice of preconditioner is $\mathbf{P}_{\mu}=\mathbf{I}$, which yields the standard conjugate gradient algorithm (CG) [35].
@@ 2.1.1 Convergence
PCG (and hence CG) satisfy a well-known convergence guarantee in terms of the condition number of $\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2}$; see e.g. [29]. [^1] Here we use that $(({\sqrt{\kappa}-1})/({\sqrt{\kappa}+1}))^{t}\leq\exp(-2t/\sqrt{\kappa})$ for all $\kappa\geq 1$ in order to obtain a simpler expression than the typical bound. This approximation is pessimistic when $\kappa\approx 1$, but it’s clear from our analysis that other bounds for CG can be used instead.
@@
Let $\mathbf{P}_{\mu}$ be any preconditioner. Then the $t$-th preconditioned-CG iterate corresponding to the preconditioner $\mathbf{P}_{\mu}$ satisfies
$\frac{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}-\mathsf{pcg}_{t}(\mu)\|_{\mathbf{A}_{\mu}}}{\|\mathbf{A}_{\mu}^{-1}\mathbf{b}\|_{\mathbf{A}_{\mu}}}\leq 2\exp\left(-\frac{2t}{\sqrt{\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})}}\right).$
Corollary 2.3 implies that if we can find $\mathbf{P}_{\mu}$ so that $\kappa(\mathbf{P}_{\mu}^{-1/2}\mathbf{A}_{\mu}\mathbf{P}_{\mu}^{-1/2})$ is small, then PCG converges rapidly. The choice of $\mathbf{P}_{\mu}$ minimizing this condition number is of course $\mathbf{P}_{\mu}=\mathbf{A}_{\mu}$, but this is not a practical preconditioner; if we knew $\mathbf{A}_{\mu}^{-1}$ we could easily compute the solution $\mathbf{A}_{\mu}^{-1}\mathbf{b}$. Thus, finding a suitable choice of $\mathbf{P}_{\mu}$, which balances improvements to the convergence of PCG with the cost of building/applying $\mathbf{P}_{\mu}$ is critical.
@@ 2.2 Deflation preconditioner
If $\mathbf{A}$ is poorly conditioned due to the presence of $r$ eigenvalues much larger than the remaining nrn-r eigenvalues, then we might hope to learn a good approximation of the top $r$ eigenvalues and “correct” this ill-conditioning. Towards this end, suppose we have a good rank-$r$ approximation $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ to $\mathbf{A}$, where $\mathbf{U}$ has orthonormal columns and $\mathbf{D}$ is diagonal. Intuitively, this low-rank approximation contains the information needed to remove the top eigenvalues of $\mathbf{A}$, thereby reducing the condition number of the preconditioned system. In particular, one can form the preconditioner
(9) $\mathbf{P}_{\mu}:=\frac{1}{\theta+\mu}\mathbf{U}(\mathbf{D}+\mu\mathbf{I})\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}),$
where $\theta>0$ is a parameter that must be chosen along with the factorization $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$.
It is not hard to verify that
(10) $\mathbf{P}_{\mu}^{-1}=(\theta+\mu)\mathbf{U}(\mathbf{D}+\mu\mathbf{I})^{-1}\mathbf{U}^{\mathsf{T}}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\mathsf{T}}).$
This means applying $\mathbf{P}_{\mu}^{-1}$ to a vector can be done in $O(dr)$ arithmetic operations, which is relatively cheap if $r\ll d$. Hence, so long as a reasonable factorization $\mathbf{U}\mathbf{D}\mathbf{U}^{\mathsf{T}}$ can be obtained, Eq. 9 can be used as a preconditioner; see [24, 33] and the references within.